(1)HOW many ways 4 letter can be selected from
the word examination .

Grade:12th pass

1 Answers

Arun
4 years ago
We have, 11 letters: {A, A, E, I, I, O, M, N, N, T, X};
Out of them there are 3 pairs: {A, A}, {I, I} and {N, N}. 
So, # of distinct letters is 8: {A, E, I, I, O, M, N, N, T, X} 

As pointed out, there are 3 different cases possible.
{abcd} all 4 letters are distinct:
8C4 = 70
 
{aabc} - two letters are alike and other two are distinct:
3C1 * 7C2 = 63
 
{aabb} - two letters are alike and other two letters are also alike:
3C2 = 3
 
Total=70+63+3=136.